OpenStudy (anonymous):

original question: solve the system of equations. 3x-5y=16 -5x+7y=-16 i multiplied the top by 5 and the bottom by 3 and got 15x-25y=80 -15x+21y=-48 Now what is x & y?

OpenStudy (anonymous):

now you can add the two equations together easily. 15x-15x-24y+21y=80-48. so -24y+21y=80-48, and -3y=32, just solve for y. then plug that value into EITHER equation (doesn't matter which) and solve for x

OpenStudy (anonymous):

*correction -25y not -24y, sorry!

OpenStudy (anonymous):

but 32 divided by -3 doesnt work out

OpenStudy (anonymous):

why not? y=32/-3

OpenStudy (anonymous):

so 15x-25(32/-3)=80

OpenStudy (anonymous):

oops, no i made a mistake it is -25y not -24y should be -4y=32

OpenStudy (anonymous):

15x-25(-8)=80

OpenStudy (anonymous):

15x-200=80

OpenStudy (anonymous):

Are these slopes?